OpenStudy (anonymous):

a fishing boat leaves its home port and travels 150 miles directly east. It then changes course and travels 40 mi due north. How long will the direct return trip take if the boat averages 23 mi/h? Solve using vectors.

OpenStudy (anonymous):

@aum Would you mind helping again? :P

OpenStudy (aum):

Position of the boat = 150i + 40j Magnitude of the vector = \(\sqrt{150^2 + 40^2} = 155.24\) miles. speed = distance / time time = distance / speed = 155.24 / 23 hours = 6.75 hours = 6 hours and 45 minutes.

OpenStudy (anonymous):

Thank you soooo much!!!

OpenStudy (aum):

You are welcome.